# Asymmetry in Spectral Graph Theory: Harmonic Analysis on Directed Networks via Biorthogonal Bases

### v1

@@ 1.2 Position relative to graph Fourier analysis
Graph Fourier analysis is often introduced through symmetric operators (undirected Laplacians/adjacencies), which guarantee orthogonal eigenvectors and stable spectral coordinates [2, 3, 16]. Directed graph settings typically replace symmetry by alternative constructions: optimization-based directed transforms and directed Laplacians associated with random walks and Cheeger-type theory [4, 1], as well as survey-level treatments of directed-graph signal processing [19]. Our approach is complementary and more “operator first”: we start from the canonical diffusion operator $P$ and develop an exact biorthogonal Fourier calculus for directed diffusion, with a transparent symmetry/asymmetry interpretation in terms of reversibility [9, 11].
@@ 4.1 Left/right eigenvectors and BGFT
Define $U^{\ast}:=V^{-1}$ so $U^{\ast}V=I$ and $u_{k}^{\ast}v_{\ell}=\delta_{k\ell}$.
@@ 4.3 Diffusion-consistent frequency ordering
Low $\omega_{\mathrm{diff}}$ corresponds to persistent/slow modes; high $\omega_{\mathrm{diff}}$ corresponds to fast decay. In the symmetric undirected limit (where $\lambda\in[-1,1]$ for normalized settings), this ordering aligns with classical low/high frequency intuition.
@@ 5.2 Operator-norm stability: diffusion and filtering
Remark 5.5 (Symmetry/asymmetry interpretation).
When $P$ is normal and diagonalizable by a unitary basis, $\operatorname{cond}(V)=1$ and the bounds become tight and symmetry-like. For non-normal $P$, $\operatorname{cond}(V)$ can be large, creating instability even if $|\lambda_{k}|\leq 1$.Sharper growth control for non-normal matrices can be expressed via pseudospectral/Kreiss-type constants; see [15, 7].

### v2

@@ 1.2 Position relative to graph Fourier analysis
Graph Fourier analysis is often introduced through symmetric operators (undirected Laplacians/adjacencies), which guarantee orthogonal eigenvectors and stable spectral coordinates [2, 3, 16]. Directed graph settings typically replace symmetry by alternative constructions: optimization-based directed transforms that seek to minimize a directed variation [4], or transforms based on the Jordan decomposition of the adjacency matrix $A$ [3]. While these approaches provide powerful tools for signal compression, they often decouple the transform from the underlying physics of diffusion. Our approach is complementary and more “operator first”: we start from the canonical diffusion operator $P$ and develop an exact biorthogonal Fourier calculus for directed diffusion, with a transparent symmetry/asymmetry interpretation in terms of reversibility [9, 11]. Unlike methods that force orthogonality through isometric embeddings, our BGFT embraces the biorthogonal geometry inherent in non-reversible Markov chains.
@@ 4.1 Left/right eigenvectors and BGFT
Remark 4.2 (Non-diagonalizable operators).
While Assumption 4.1 holds for almost all transition matrices (diagonalizable matrices are dense in $\mathbb{C}^{n\times n}$), certain highly symmetric digraph structures can yield non-diagonalizable $P$. In such cases, the BGFT can be generalized using the Jordan canonical form or the Schur decomposition. However, the biorthogonal framework presented here focuses on the most common case where a complete set of eigenvectors exists, providing a direct physical link to decay modes.
@@ 4.3 Diffusion-consistent frequency ordering
Low $\omega_{\mathrm{diff}}$ corresponds to persistent/slow modes; high $\omega_{\mathrm{diff}}$ corresponds to fast decay. Compared to the imaginary-part ordering (phase) often used in adjacency-based GFTs, $\Re(1-\lambda)$ provides a direct measure of spectral energy dissipation. This choice aligns with the variational property of the random-walk Laplacian, where $\Re(1-\lambda)$ quantifies the smoothness of a mode relative to the one-step diffusion process.
@@ 5.2 Operator-norm stability: diffusion and filtering
Remark 5.5 (Symmetry/asymmetry and Instability).
When $P$ is normal and diagonalizable by a unitary basis, $\operatorname{cond}(V)=1$ and the bounds become tight and symmetry-like. It is critical to distinguish between structural asymmetry ($\alpha(P)>0$) and numerical instability ($\operatorname{cond}(V)\gg 1$). Asymmetry is a prerequisite for the loss of orthogonality, but it does not necessitate instability; for instance, the directed cycle is maximally asymmetric but remains normal and perfectly stable. For non-normal $P$, $\operatorname{cond}(V)$ can be large, creating instability even if $|\lambda_{k}|\leq 1$. Our numerical experiments in Section 9 confirm that this ill-conditioning, rather than directedness per se, drives the reconstruction error.
@@ 9.2 Tasks and metrics
  • Perturbed directed cycle $\overrightarrow{C_{n}}^{(\varepsilon)}$: add a directed chord from node 0 to node $n/2$ with weight $\varepsilon=20$, then renormalize rows to keep $P$ stochastic; this typically yields non-normal $P$ and large $\operatorname{cond}(V)$.